# Hallucination Detection in LLMs: Fast and Memory-Efficient Finetuned Models

### v1

@@ Abstract
Uncertainty estimation is a necessary component when implementing AI in high-risk settings, such as autonomous cars, medicine, or insurances. Large Language Models (LLMs) have seen a surge in popularity in recent years, but they are subject to hallucinations, which may cause serious harm in high-risk settings. Despite their success, LLMs are expensive to train and run: they need a large amount of computations and memory, preventing the use of ensembling methods in practice. In this work, we present a novel method that allows for fast and memory-friendly training of LLM ensembles. We show that the resulting ensembles can detect hallucinations and are a viable approach in practice as only one GPU is needed for training and inference. Code will be made available upon acceptance.
Figure 1: (Left) The ensemble utilizes a shared matrix of pre-trained “slow weights,” which are updated with LoRA matrices during training and then merged. Each ensemble member is represented by an individual rank-one matrix (fast weights) that is combined with the shared weights using a Hadamard product. (Right) The ensemble generates uncertainty estimates, which serve as features for a classifier to determine whether the LLM’s prediction is correct or hallucinated.
@@ 1 Introduction
To address these limitations, we propose a fast and memory-efficient deep ensemble method which is able to provide reliable uncertainty estimates. Figure 1 describes our proposed method, where low-rank matrices are added on top of a pretrained model and used for fine-tuning. A Low-Rank Adaptation (LoRA) matrix allows the whole ensemble to be trained cheaply and flexibly [hu2021lora], and each ensemble member has an individual rank-one fast weight matrix [^1] LoRA matrices are rank-$r$ matrices added elementwise to the weight matrix, while the fast weights – rank-1 matrices – are multiplied elementwise.. This allows the members to be diverse. An uncertainty metric is used afterward and fed into a binary classifier, which is trained to discriminate between hallucinated and correct predictions.
@@ 2 Related Work
Recent studies have proposed a memory-friendly strategy that fine-tunes LLM ensembles from pre-trained weights, rather than training from scratch  [wang_lora_2023, zhai2023uncertaintyloraensemble]. This method, referred to as a LoRA Ensemble, assigns each ensemble member its own set of LoRA matrices  [hulora]. While this approach has been utilized to compute uncertainty estimates  [wang_lora_2023, balabanovloraensembles, zhai2023uncertaintyloraensemble], it has not been specifically applied to hallucination detection tasks. Our approach is similar in two ways: it reduces training costs by utilizing pre-trained weights and it employs LoRA matrices during fine-tuning. However, unlike the LoRA Ensemble, our method does not rely on LoRA matrices to represent ensemble members. Instead, after fine-tuning, we merge the LoRA matrices with the pre-trained weights and represent the ensemble using sets of rank-one fast weights.
@@ 3 Method
$\begin{split}&\mathcal{H}\left[P(\mathbf{x}_{t}|\mathbf{x}_{<t};\mathcal{D})\right]=\\ &-\sum_{x_{t}}P(x_{t}|\mathbf{x}_{<t};\mathcal{D})\log P(x_{t}|\mathbf{x}_{<t};\mathcal{D}).\end{split}$ (1)
The predictive entropy can further be divided into its two subcomponents aleatoric and epistemic uncertainties:
@@ 4 Design of experiments
Baselines
We evaluate our proposed adaptations to BatchEnsemble during fine-tuning, including BatchEnsemble with noise injection (NI). For uncertainty-based experiments, we compare against two baselines: a prompt-based method that approximates the output distribution using repeated prompting with stochastic sampling (temperature = 0.5, top-p = 0.99, top-k = 5)  [kuhnsemanticuncertainty, chen_inside_2024], and a LoRA Ensemble, which applies regularization to LoRA B matrices as described by Wang et al. on MMLU and SQuAD datasets  [wang_lora_2023]. All models, including baselines, are fine-tuned on all modules and utilize the uncertainty measurements described in Equation (LABEL:uncertainty_measures) for consistency. We also evaluate our methods against a single model fine-tuned from pre-trained weights to ensure fair performance comparison. All models utilize Mistral-7B-Instruct-v0.2 as pre-trained weights  [jiangmistral].
@@ 5 Results
In Table 1, we assess the quality of our models’ and baselines’ uncertainty estimates by using them as features for classifiers to predict hallucinations and OOD instances. The uncertainty estimates are fed into five classifiers: k-Nearest Neighbors, logistic regression, decision trees, random forest, and support vector machine. We report the top-1 classification accuracy for each model’s uncertainty estimates. See Appendix for details.
@@
BatchEnsemble 97.8 68.0 62.4
BatchEnsemble+NI 96.5 66.9 61.9
@@
BatchEnsemble 85.9 93.4 56.7
BatchEnsemble+NI 85.4 92.6 53.2
@@ 6 Conclusion
As LLMs are increasingly used across various fields, minimizing harmful predictions is crucial. This work suggests that deep ensembles are effective for quantifying uncertainty and detecting hallucinations.
Our results indicate that while current uncertainty estimation techniques are versatile, further development is needed to enhance their effectiveness in detecting factual hallucinations and distribution shifts. Nevertheless, uncertainty estimation shows promise for identifying faithful hallucinations. Integrating uncertainty estimates with classifiers could help mitigate instances where LLMs deviate from instructions. Future research could explore combining this approach with strategies such as prompting users for clarification when instructions are ambiguous.
While we used entropy-based metrics to measure uncertainty, other methods such as EigenScore  [chen_inside_2024] or semantic entropy  [kuhnsemanticuncertainty] might offer better estimates for specific tasks. BatchEnsemble could incorporate these metrics, potentially reducing inference time by requiring just one forward pass.
Finally, we hypothesized that aleatoric and epistemic uncertainty may correspond to faithfulness and factual hallucinations, respectively. Our findings suggest that current methods for measuring epistemic uncertainty are insufficiently diverse to confirm this hypothesis. However, future noise injection methods may open the path to more informative uncertainty estimates while maintaining predictive performance.
@@ Acknowledgments
This project was funded by The Kjell and Märta Beijer Foundation, WASP, ERC grant 101054643. The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725.

### v2

# Hallucination Detection in LLMs: Fast and Memory-Efficient Fine-Tuned Models

@@ Abstract
Uncertainty estimation is a necessary component when implementing AI in high-risk settings, such as autonomous cars, medicine, or insurances. Large Language Models (LLMs) have seen a surge in popularity in recent years, but they are subject to hallucinations, which may cause serious harm in high-risk settings. Despite their success, LLMs are expensive to train and run: they need large amounts of compute and memory, preventing the use of ensembling methods in practice. In this work, we present a novel method that allows for fast and memory-friendly training of LLM ensembles. We show that the resulting ensembles can detect hallucinations and are a viable approach in practice as only one GPU is needed for training and inference. Code available at: https://github.com/Gabriel-Arteaga/LLM-Ensemble
Figure 1: (Left) The ensemble utilizes a shared matrix of pre-trained “slow weights” $U$, which are updated with LoRA matrices ($BA$) during training and then merged. Each ensemble member is represented by an individual rank-one matrix (fast weights $V$) that is combined with the shared weights using a Hadamard product. (Right) The ensemble generates uncertainty estimates, which serve as features for a classifier to determine whether the LLM’s prediction is correct or hallucinated.
@@ 1 Introduction
To address these limitations, we propose a fast and memory-efficient deep ensemble method which is able to provide reliable uncertainty estimates. Figure 1 illustrates our proposed method, where Low-Rank Adaptation (LoRA) matrices are added on top of a pretrained model and used for fine-tuning. LoRA allows the whole ensemble to be trained cheaply and flexibly [hu2021lora], with each ensemble member having an individual rank-one fast weight matrix [^1] LoRA matrices are rank-$r$ matrices added elementwise to the weight matrix, while the fast weights – rank-1 matrices – are multiplied elementwise.. This allows the members to be diverse. An uncertainty metric is used afterward and fed into a binary classifier, which is trained to discriminate between hallucinated and correct predictions.
@@ 2 Related Work
Recent studies have proposed a memory-friendly strategy that fine-tunes LLM ensembles from pre-trained weights, rather than training from scratch  [wang_lora_2023, zhai2023uncertaintyloraensemble]. This method, referred to as a LoRA Ensemble, assigns each ensemble member its own set of LoRA matrices  [hulora]. While this approach has been utilized to compute uncertainty estimates  [wang_lora_2023, balabanovloraensembles, zhai2023uncertaintyloraensemble], it has not been specifically applied to hallucination detection tasks. Our approach is similar in two ways: it reduces training costs by utilizing pre-trained weights, and it employs LoRA matrices during fine-tuning. However, unlike the LoRA Ensemble, our method does not rely on LoRA matrices to represent ensemble members. Instead, after fine-tuning, we merge the LoRA matrices with the pre-trained weights and represent the ensemble using sets of rank-one fast weights.
@@ 3 Method
$\begin{split}&\mathcal{H}\left[P(\mathbf{x}_{t}|\mathbf{x}_{<t};\mathcal{D})\right]=\\ &-\sum_{x_{t}}P(x_{t}|\mathbf{x}_{<t};\mathcal{D})\log P(x_{t}|\mathbf{x}_{<t};\mathcal{D})\end{split}$ (1)
where $\mathcal{D}$ refers to the overall training data distribution. The predictive entropy can further be divided into its two subcomponents aleatoric and epistemic uncertainties:
@@ 4 Design of experiments
Models and Baselines
We evaluate our proposed adaptations to BatchEnsemble during fine-tuning, including BatchEnsemble with noise injection (NI). For uncertainty-based experiments, we compare against two baselines: a sample-based method that approximates the output distribution using repeated prompting with stochastic sampling (temperature = 0.5, top-p = 0.99, top-k = 5)  [kuhnsemanticuncertainty, chen_inside_2024], and a LoRA Ensemble, which applies regularization to LoRA B matrices as described by Wang et al. on MMLU and SQuAD datasets  [wang_lora_2023].
All models, including baselines, are fine-tuned on all modules using LoRA adapters with a rank of $r=8$ and a scaling factor of $\alpha=32$. For BatchEnsemble, the LoRA matrices are merged into the shared weights after training. Unless otherwise specified, we use an ensemble size of 4 across all experiments. This ensemble size determines the number of rank-1 matrices for BatchEnsemble, the number of adapters for the LoRA ensemble, and the number of samples generated by the sample-based method. Additionally, all models, including baselines, use the uncertainty measures described in Equation (LABEL:uncertainty_measures).
We also evaluate our methods against a single model fine-tuned from pre-trained weights to ensure fair performance comparison. All models in this study leverage Mistral-7B-Instruct-v0.2 as a core component, either by directly fine-tuning the pre-trained weights or by incorporating them into more complex architectures like BatchEnsemble, where the shared weights are replaced by the pre-trained weights [jiangmistral].
@@
Classifier Training Details
To evaluate the quality of the uncertainty estimates produced by our method and the baseline approaches, we trained five distinct classifiers: k-Nearest Neighbors, logistic regression, decision tree, random forest, and support vector machine. These classifiers were trained to distinguish between hallucinated and non-hallucinated predictions using uncertainty estimates derived from the outputs of our method and the baselines.
For each task—faithfulness hallucination detection, factual hallucination detection, and Out-of-Distribution (OOD) detection—we created custom datasets based on the outputs of our method and the baseline approaches. In all cases, 5000 data points were propagated through the models to generate predictions, which were then annotated as hallucinated or non-hallucinated depending on the task-specific criteria. The resulting datasets were split into 80% for training and 20% for testing, with the split performed randomly using a seed of 42 to ensure consistency across experiments.
The annotation methods for hallucinations were task-specific. For faithfulness hallucination detection, as described earlier, any response to unanswerable questions in the SQuAD v2 dataset that deviated from ”I don’t know” was labeled as hallucinated. For factual hallucination detection, we used multiple-choice questions from the MMLU dataset, where incorrect answers were labeled as hallucinated and correct answers as non-hallucinated. For OOD detection, we followed the same annotation procedure as faithfulness hallucination detection, using the same set of unanswerable questions from SQuAD v2, with the primary difference being how our method and the baselines were trained for this task.
@@ 5 Results
In Table 1, we report the top-1 classification accuracy achieved by the best-performing classifier on the uncertainty estimates from our baselines and models. Detailed classification accuracies for each classifier are provided in Appendix B.
@@
(Ours) BatchEnsemble 97.8 68.0 62.4
(Ours) BatchEnsemble+NI 96.5 66.9 61.9
@@
(Ours) BatchEnsemble 85.9 93.4 56.7
(Ours) BatchEnsemble+NI 85.4 92.6 53.2
@@
6 Discussion
This paper focused on exploring uncertainty-based methods for detecting hallucinations in LLMs, and demonstrating the possibility of faithfulness and factual hallucination detection. While comparing against methods beyond uncertainty estimation could provide valuable insights, no standardized benchmark currently exists for systematic comparison between uncertainty-based and other detection strategies. We framed hallucination detection as a binary classification task and designed benchmarks accordingly. However, creating a comprehensive benchmark that incorporates other methods meaningfully remains a challenge. Addressing this gap presents a promising direction for future research.
Another area for future exploration is the use of alternative metrics beyond predictive entropy and its subcomponents. Methods such as EigenScore [chen_inside_2024] or semantic entropy [kuhnsemanticuncertainty] may offer better performance for specific tasks by capturing uncertainty in more nuanced ways. Integrating these metrics into approaches like BatchEnsemble could enhance both efficiency and robustness of hallucination detection by reducing inference time while maintaining or improving reliability.
Our experiments were limited to exploring the method with a single pretrained language model. However, we hypothesize that the method works with other LLM sizes and architectures. Future work could explore the impact of factors such as model size on overconfidence and whether this is reflected in the model’s uncertainty estimates, providing further insights into the robustness and scalability of uncertainty-based hallucination detection.
Finally, we believe that aleatoric and epistemic uncertainty may correspond to faithfulness and factual hallucinations, respectively. Our findings suggest that current methods for measuring epistemic uncertainty are not diverse enough to confirm this hypothesis. Future work could involve developing more sophisticated noise injection techniques to generate diverse ensemble members, which may enhance the accuracy of epistemic uncertainty measurements while maintaining predictive performance. Such efforts could open new research directions to better understand how different uncertainty components relate to faithfulness and factual hallucinations.
@@ 7 Conclusion
We have developed a memory-efficient method to fine-tune LLMS using LoRA matrices and rank-1 matrix modifications. A key contribution of our work is demonstrating how pretrained models can be enhanced with an uncertainty estimation component to enable effective hallucination detection. Our approach demonstrated the ability to distinguish between hallucinated and non-hallucinated content, achieving high accuracy in faithfulness detection, even in resource-constrained settings.
While our uncertainty-based methods provide valuable tools for hallucination detection, and have shown strong performance in detecting faithfulness hallucinations, challenges remain, particularly in detecting factual hallucinations and managing out-of-distribution data points. Nevertheless, our approach is an important step towards reducing harmful outputs from LLMs, which is crucial for deploying these models in safety-critical environments.
The combination of uncertainty estimation with flexible fine-tuning approaches like BatchEnsemble and LoRA shows strong potential for enhancing LLM reliability and robustness. Future work could build on these findings by integrating more sophisticated uncertainty metrics and establishing standardized benchmarks, paving the way for more effective and generalizable hallucination detection strategies.
@@ Acknowledgments
This work builds upon research originally conducted as part of Gabriel Y. Arteaga’s master thesis at Uppsala University [arteaga2024hallucination].
This research was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by Knut and Alice Wallenberg Foundation, by Kjell och Märta Beijer Foundation, and ERC grant 101054643. The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725.